Ushna has enough money to buy 24 apples. If the price of each apple is reduced by 5 cents, she will be able to buy an extra 6 apples with the same sum of money. Find the original cost of the apple.
*Prove your answer is 25 cents*

Answer by mananth(16949)

Ushna has enough money to buy 24 apples. If the price of each apple is reduced by 5 cents, she will be able to buy an extra 6 apples with the same sum of money. Find the original cost of the apple.
let apple cost x cents
..
she has 24x cents
..
(x-5) cost of apple
24x/(x-5)= 30
multiply by (x-5)
24x=(x-5)*30
24x=30x-150
-30 x
24x-30x=-150
-6x=-150
/-6
x=25 cents ------original cost of apple
...
m.ananth@hotmail.ca